### v1

# An Ecosystem for Ontology Interoperability [Vision]

Abstract.
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications. A case study in the building domain validates the usefulness of the proposed ecosystem.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-interoperability.
1. Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, visit, Sydney) alone, but with a logic-based expression (Person, TravelAction, Place) derived from the schema.org vocabulary (Schema.org, 2025), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited. The triple (Sydney, visit, Sydney) describes an event in which the American actress Sydney Sweeney visited the city of Sydney (Mitchell, 2023).
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain. There is no guarantee that ontologies will agree on the same conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies. For example, it is common for ontologies to encompass a range of different concepts. A person can be described using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee). Moreover, the terminology used in ontologies can be different. It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
2. Related Work
2.1. Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies (Gangemi and Presutti, 2009). Figure 1 illustrates that ODPs use the star-based matching schema. Once the ODP is determined, each new ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of the ontology.
Figure 1. Ontology design patterns (ODPs).
[…]
Ontology design patterns (ODPs).
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies. They are static and are not prone to capturing changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
2.2. Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies (Euzenat and Shvaiko, 2013) or different versions of the same ontology (Klein and Fensel, 2001). Figure 2 shows that OM&OV conform to the pairwise matching schema. This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
Figure 2. Ontology matching and versioning (OM&OV).
[…]
Ontology matching and versioning (OM&OV).
OM&OV are typically resource-consuming. The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
2.3. Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) (Qiang, 2023). Figure 3 shows that the OCKGs comply with the data-driven schema. While ontologies comply with KG(s), their reference mappings can be derived from similar concepts pointing to the same data point.
Figure 3. Ontology-compliant knowledge graphs (OCKGs).
[…]
Ontology-compliant knowledge graphs (OCKGs).
Building OCKGs is not easy. This process requires extensive workloads to find the alignments between different ontologies, so that each KG can be modelled with different ontologies to determine which ontology fits the KG(s) better.
3. Our Vision: An Ecosystem
Figure 4 illustrates an overview of the ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
Figure 4. An overview of the ecosystem.
[…]
An overview of the ecosystem.
The design phase is responsible for understanding the requirements and creating a blueprint or prototype. In ontology engineering, these requirements are collected by competency questions. ODPs can serve as a skeleton to create conceptual models based on competency questions. The develop phase includes turning the prototype into a full product and applying a unit test. In ontology engineering, this phase is referred to as ontology construction and evaluation. OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version. The deploy phase usually refers to implementing the product in the real world. In ontology engineering, we refer to this as applying the ontology to actual KGs. We can utilise OCKGs to comprehend the interoperability of ontologies within each KG or a group of KGs, thereby enhancing the design of the ontology in the next cycle.
4. A Case Study in the Building Domain
Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades. While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain (Qiang et al., 2023). We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability. The running example is described as follows: A sensor is installed in the room to monitor the room temperature. The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
4.1. Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is recommended to build an ontology based on ODPs. Several ODPs can be used in this case (shown in Figure 5).
  1. (1)
    Observation Pattern (used in SOSA/SSN (Haller et al., 2017)). Sensor observation is defined as a type of observation. The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
  2. (2)
    Measurement Pattern (used in the early version of SAREF (Daniele et al., 2015)). Sensor observation is defined as a measurement. The measurement is made by the sensor, which is a measure of the room related to the property temperature.
  3. (3)
    Event Pattern (used in DUL (Presutti and Gangemi, 2016)). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
  4. (4)
    Activity Pattern (used in PROV-O (Lebo et al., 2013)). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
Figure 5. Modelling using different ODPs.
[…]
Modelling using different ODPs.
While all of these ODPs can effectively capture the main concepts in the example, they have differences that need to be considered.
  1. (1)
    The definition of the concept can be different. For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
  2. (2)
    The levels of expressiveness are different. Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish whether it is a structural attachment (i.e. isPartOf) or only a spatial presence (i.e. isLocatedIn).
4.2. Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN. As illustrated in Figure 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. If we consider SOSA to be an evolved version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV tasks can be unified with minor modifications. OM&OV tasks can be classified into two main categories.
Figure 6. Aligning SOSA and SSN.
[…]
Aligning SOSA and SSN.
  1. (1)
    Finding trivial alignments only considers the syntactic matching of the entity. For example, the mappings between the classes sosa:Sensor and ssn:Sensor, and the properties sosa:madeBySensor and ssn:madeBySensor. We also refer to these mappings as “seed mappings” because they serve as the basis for subsequent lexical and semantic matching.
  2. (2)
    Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity. For example, the mappings between the classes sosa:ObservableProperty and ssn:Property, and the properties sosa:isHostedBy and ssn:onPlatform. The alignment should also exclude non-existent and counter-intuitive mappings. For example, ssn:hasProperty should not have a direct SOSA equivalent mapping.
4.3. Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it remains challenging to determine which ontology to use in the real world. Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation. Figure 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
  1. (1)
    The stimulus “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
  2. (2)
    The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings.
  1. (1)
    We use ssn:Property to replace sosa:ObservableProperty because this class is more general and can be applied in a broader context for sensing, including actuation and sampling.
  2. (2)
    We use sosa:isHostedby to replace ssn:onPlatform because this property is more closely related to describing a sensor physically installed in the room.
  3. (3)
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish these two concepts to avoid confusion.
Figure 7. Build OCKGs using SOSA and SSN.
[…]
Building OCKGs using SOSA and SSN.
Figure 8 shows the evolved ontology and the OCKG built on the evolved ontology. Apparently, the evolved ontology is more suitable for the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
Figure 8. Evolved ontology and new OCKG.
[…]
Evolved ontology and new OCKG.
4.4. Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align the different concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
5. Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Introduced in 2023, ontology-compliant KGs aim to explore the interoperability between ontology and its complementary KGs.
While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and constructing OCKGs in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
Acknowledgements.
The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v2

@@
-An Ecosystem for Ontology Interoperability [Vision]
-Abstract.
+An Ecosystem for Ontology Interoperability
+Abstract
 Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs).
@@
 The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications.
-A case study in the building domain validates the usefulness of the proposed ecosystem.
-PVLDB Reference Format:
-PVLDB, 19(1): XXX-XXX, 2025.
-doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
-doi:$X$.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-interoperability.
-1. Introduction
+A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
+Index Terms:
+ontology, data interoperability
+I Introduction
 For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs.
 While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs.
-For example, it can be challenging to understand a KG snippet (Sydney, visit, Sydney) alone, but with a logic-based expression (Person, TravelAction, Place) derived from the schema.org vocabulary (Schema.org, 2025), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited.
-The triple (Sydney, visit, Sydney) describes an event in which the American actress Sydney Sweeney visited the city of Sydney (Mitchell, 2023).
+For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a logic-based expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited.
+The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
 Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain.
@@
 We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability.
-We propose an ecosystem for ontology interoperability and use a case study in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
-2. Related Work
-2.1. Ontology Design Patterns
-The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies (Gangemi and Presutti, 2009).
-Figure 1 illustrates that ODPs use the star-based matching schema.
-Once the ODP is determined, each new ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of the ontology.
-Figure 1. Ontology design patterns (ODPs).
-[…]
-Ontology design patterns (ODPs).
+We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
+II Related Work
+II-A Ontology Design Patterns
+The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2].
+Fig.
+1 illustrates that ODPs use the star-based matching schema.
+Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
+Figure 1: Ontology design patterns (ODPs).
 However, ODPs also have drawbacks.
@@
 Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
-2.2. Ontology Matching and Versioning
-Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies (Euzenat and Shvaiko, 2013) or different versions of the same ontology (Klein and Fensel, 2001).
-Figure 2 shows that OM&OV conform to the pairwise matching schema.
+II-B Ontology Matching and Versioning
+Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4].
+Fig.
+2 shows that OM&OV conform to the pairwise matching schema.
 This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
-Figure 2. Ontology matching and versioning (OM&OV).
-[…]
-Ontology matching and versioning (OM&OV).
+Figure 2: Ontology matching and versioning (OM&OV).
 OM&OV are typically resource-consuming.
@@
 TBox-to-ABox matching).
-2.3. Ontology-Compliant Knowledge Graphs
-Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) (Qiang, 2023).
-Figure 3 shows that the OCKGs comply with the data-driven schema.
-While ontologies comply with KG(s), their reference mappings can be derived from similar concepts pointing to the same data point.
-Figure 3. Ontology-compliant knowledge graphs (OCKGs).
-[…]
-Ontology-compliant knowledge graphs (OCKGs).
+II-C Ontology-Compliant Knowledge Graphs
+Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) [5].
+Fig.
+3 shows that the OCKGs comply with the star-based matching schema.
+While ontologies comply with KG(s), their reference mappings can be derived from the concepts pointing to the same data instance.
+Figure 3: Ontology-compliant knowledge graphs (OCKGs).
 Building OCKGs is not easy.
 This process requires extensive workloads to find the alignments between different ontologies, so that each KG can be modelled with different ontologies to determine which ontology fits the KG(s) better.
-3. Our Vision: An Ecosystem
-Figure 4 illustrates an overview of the ecosystem for ontology interoperability.
+III An Ecosystem
+Fig.
+4 illustrates an overview of the ecosystem for ontology interoperability.
 Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy.
 We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
-Figure 4. An overview of the ecosystem.
-[…]
-An overview of the ecosystem.
+Figure 4: An overview of the ecosystem.
 The design phase is responsible for understanding the requirements and creating a blueprint or prototype.
@@
 We can utilise OCKGs to comprehend the interoperability of ontologies within each KG or a group of KGs, thereby enhancing the design of the ontology in the next cycle.
-4. A Case Study in the Building Domain
+IV A Case Study in the Building Domain
 Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades.
-While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain (Qiang et al., 2023).
+While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain [6].
 We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability.
@@
 The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
-4.1. Design Phase
+IV-A Design Phase
 Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability.
 It is recommended to build an ontology based on ODPs.
-Several ODPs can be used in this case (shown in Figure 5).
+Several ODPs can be used in this case, as illustrated in Fig.
+5.
   1. (1)
-    Observation Pattern (used in SOSA/SSN (Haller et al., 2017)).
+    Observation Pattern (used in SOSA [7] and SSN [8]).
 Sensor observation is defined as a type of observation.
@@
   2. (2)
-    Measurement Pattern (used in the early version of SAREF (Daniele et al., 2015)).
+    Measurement Pattern (used in the early version of SAREF [9]).
 Sensor observation is defined as a measurement.
@@
   3. (3)
-    Event Pattern (used in DUL (Presutti and Gangemi, 2016)). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
+    Event Pattern (used in DUL [10]). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
   4. (4)
-    Activity Pattern (used in PROV-O (Lebo et al., 2013)).
-Sensor observation is defined as an activity.
-The activity is associated with the sensor that uses the room and generates a temperature measurement.
-Figure 5. Modelling using different ODPs.
-[…]
-Modelling using different ODPs.
+    Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
 While all of these ODPs can effectively capture the main concepts in the example, they have differences that need to be considered.
-  1. (1)
+  1. 1.
     The definition of the concept can be different.
@@
 Activity Pattern defines the temperature as a member of the room.
-  2. (2)
+  2. 2.
     The levels of expressiveness are different.
@@
 isLocatedIn).
-4.2. Develop Phase
+Figure 5: Modelling using different ODPs.
+IV-B Develop Phase
 We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN.
-As illustrated in Figure 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs.
+As illustrated in Fig.
+6, we can see that SOSA and SSN are very similar because they are built from the same ODPs.
 However, they are not exactly the same.
 The OM task is to find aligned concepts between two different ontologies.
-If we consider SOSA to be an evolved version of SSN, then this OM task is also an OV task.
-We use OM&OV here because OM and OV tasks can be unified with minor modifications.
+If we consider SOSA to be an updated version of SSN, then this OM task is also an OV task.
+We use OM&OV here because OM and OV can be unified with minor modifications [12].
 OM&OV tasks can be classified into two main categories.
-Figure 6. Aligning SOSA and SSN.
-[…]
-Aligning SOSA and SSN.
-  1. (1)
+Figure 6: Aligning SOSA and SSN.
+  1. 1.
     Finding trivial alignments only considers the syntactic matching of the entity.
@@
 We also refer to these mappings as “seed mappings” because they serve as the basis for subsequent lexical and semantic matching.
-  2. (2)
+  2. 2.
     Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity.
@@
 The alignment should also exclude non-existent and counter-intuitive mappings.
-For example, ssn:hasProperty should not have a direct SOSA equivalent mapping.
-4.3. Deploy Phase
+For example, ssn:hasProperty should not have a direct equivalent mapping in SOSA.
+IV-C Deploy Phase
 Now we have developed SOSA and SSN, along with their reference alignment.
@@
 Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation.
-Figure 7 shows the OCKGs based on SOSA and SSN.
+Fig.
+7 shows the OCKGs based on SOSA and SSN.
 We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
-  1. (1)
-    The stimulus “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
-  2. (2)
+  1. 1.
+    The stimulus of the sensor “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
+  2. 2.
     The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
 To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA.
-We also make several design choices between aligned mappings.
-  1. (1)
+We also make several design choices between aligned mappings (retrieved from [13, 14, 15]).
+  1. 1.
     We use ssn:Property to replace sosa:ObservableProperty because this class is more general and can be applied in a broader context for sensing, including actuation and sampling.
-  2. (2)
+  2. 2.
     We use sosa:isHostedby to replace ssn:onPlatform because this property is more closely related to describing a sensor physically installed in the room.
-  3. (3)
+  3. 3.
     We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest.
-Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish these two concepts to avoid confusion.
-Figure 7. Build OCKGs using SOSA and SSN.
-[…]
-Building OCKGs using SOSA and SSN.
-Figure 8 shows the evolved ontology and the OCKG built on the evolved ontology.
-Apparently, the evolved ontology is more suitable for the provided KG instance.
+Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
+Figure 7: Build OCKGs using SOSA and SSN.
+Fig.
+8 shows the updated ontology and the OCKG built on the updated ontology.
+Apparently, the updated ontology is more suitable for the provided KG instance.
 In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times.
@@
 For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
-Figure 8. Evolved ontology and new OCKG.
-[…]
-Evolved ontology and new OCKG.
-4.4. Implications
+Figure 8: Updated ontology and new OCKG.
+IV-D Implications
 Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks.
-As a complement to ODPs, OM&OV can align the different concepts that are not captured by ODPs.
+As a complement to ODPs, OM&OV can align the concepts that are not captured by ODPs.
 On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems.
 OM&OV is a critical middleware between ODPs and OCKGs.
-5. Discussion & Conclusion
+V Discussion & Conclusion
 Ontology interoperability has been a long-standing issue studied in the Semantic Web community.
@@
 We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and constructing OCKGs in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
-Acknowledgements.
+Supplemental Material Statement
+The source code, data, and/or other artifacts used in this study have been made available at https://github.com/qzc438-research/ontology-interoperability.
+Acknowledgment
 The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v3

@@
 We propose an ecosystem for ontology interoperability.
-The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications.
+The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications.
 A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
@@
 While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs.
-For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a logic-based expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited.
+For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a class schema expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a person, while the object “Sydney” refers to the place being visited.
 The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
-Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain.
-There is no guarantee that ontologies will agree on the same conceptualisation.
+Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application.
+There is no guarantee that ontologies in the same domain agree on the conceptualisation.
 Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies.
-For example, it is common for ontologies to encompass a range of different concepts.
-A person can be described using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee).
+For example, it is common for ontologies to encompass a range of different concepts, such as describing a person using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee).
 Moreover, the terminology used in ontologies can be different.
 It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
-In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs).
+In this paper we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs).
 We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability.
@@
 Fig.
-1 illustrates that ODPs use the star-based matching schema.
+1 illustrates that ODPs use a star-based matching schema.
 Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
@@
 Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies.
-They are static and are not prone to capturing changes over time.
+They are static and tend not to capture changes over time.
 Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
@@
 Fig.
-2 shows that OM&OV conform to the pairwise matching schema.
-This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
+2 shows that OM&OV conform to a pairwise matching schema.
+This means that each mapping is a pairwise process to find the correspondences between two ontologies.
 Figure 2: Ontology matching and versioning (OM&OV).
-OM&OV are typically resource-consuming.
+OM&OV are typically resource-intensive.
 The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies.
@@
 II-C Ontology-Compliant Knowledge Graphs
-Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) [5].
+Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5].
 Fig.
-3 shows that the OCKGs comply with the star-based matching schema.
-While ontologies comply with KG(s), their reference mappings can be derived from the concepts pointing to the same data instance.
+3 shows that OCKGs comply with a star-based matching schema.
+Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
 Figure 3: Ontology-compliant knowledge graphs (OCKGs).
 Building OCKGs is not easy.
-This process requires extensive workloads to find the alignments between different ontologies, so that each KG can be modelled with different ontologies to determine which ontology fits the KG(s) better.
+To determine which candidate ontology fits the KG(s) better, the KG needs to be modelled with each candidate.
+This process itself is essentially building an alignment of the candidate ontologies, informed by the KG under study.
 III An Ecosystem
 Fig.
-4 illustrates an overview of the ecosystem for ontology interoperability.
+4 illustrates an overview of an ecosystem for ontology interoperability.
 Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy.
-We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
+We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve interoperability.
 Figure 4: An overview of the ecosystem.
@@
 Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability.
-It is recommended to build an ontology based on ODPs.
-Several ODPs can be used in this case, as illustrated in Fig.
-5.
+It is commonly recommended to reuse pre-existing ODPs.
+Fig.
+5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
+While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown in the following.
   1. (1)
@@
     Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
-While all of these ODPs can effectively capture the main concepts in the example, they have differences that need to be considered.
+On the other hand, differences in ontology characteristics, such as those listed below, need to be addressed in practice.
   1. 1.
-    The definition of the concept can be different.
+    The definition of the concepts can be different.
 For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room.
@@
 IV-B Develop Phase
-We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN.
+We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]).
 As illustrated in Fig.
@@
 The OM task is to find aligned concepts between two different ontologies.
-If we consider SOSA to be an updated version of SSN, then this OM task is also an OV task.
+SOSA and an updated SSN were co-developed, with a published alignment between them.
+Sosa was more of a module of the updated SSN (for which there were other modules too).
+If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task.
 We use OM&OV here because OM and OV can be unified with minor modifications [14].
@@
 Now we have developed SOSA and SSN, along with their reference alignment.
-However, it remains challenging to determine which ontology to use in the real world.
-Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation.
+However, it is hard to determine which ontology to use without applying the real-world KGs.
+Some concepts used in a KG may not have the corresponding classes and properties in the ontology.
+Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation.
 Fig.
 7 shows the OCKGs based on SOSA and SSN.
-We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
+We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts within the given KG.
   1. 1.
@@
     The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
-To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA.
+To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA.
 We also make several design choices between aligned mappings (retrieved from [13, 15, 16]).
@@
 Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings.
-Introduced in 2023, ontology-compliant KGs aim to explore the interoperability between ontology and its complementary KGs.
+Ontology-compliant knowledge graphs (OCKGs) aim to explore the interoperability between an ontology and its complementary KGs.
 While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability.
@@
 Supplemental Material Statement
-The source code, data, and/or other artifacts used in this study have been made available at https://github.com/qzc438-research/ontology-interoperability.
+The source code, data, and/or other artifacts used in this study are available at https://github.com/qzc438-research/ontology-interoperability.
 Acknowledgment
-The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+The author is supervised by Kerry Taylor (Australian National University), Weiqing Wang (Monash University), Jing Jiang (Australian National University), Subbu Sethuvenkatraman (Commonwealth Scientific and Industrial Research Organisation), Qing Wang (Australian National University), and Alex Potanin (Australian National University).
+The author thanks the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v5

@@
 We propose an ecosystem for ontology interoperability.
-The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications.
+The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and data-driven ontology validation (DOVA) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications.
 A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
-Index Terms:
-ontology, data interoperability
 I Introduction
@@
 It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
-In this paper we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs).
+In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs).
 We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability.
@@
 II-A Ontology Design Patterns
-The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2].
+The goal of ontology design patterns (ODPs) is to identify common ontology fragments across different ontologies [2].
 Fig.
@@
 This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices.
-Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies.
+Secondly, ODPs are commonly handcrafted on the basis of existing ontologies.
 They are static and tend not to capture changes over time.
@@
 OM&OV are typically resource-intensive.
-The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies.
+The complexity of pairwise mapping is the product of the number of entities in the two ontologies.
 The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist.
@@
 TBox-to-ABox matching).
-II-C Ontology-Compliant Knowledge Graphs
-Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5].
+II-C Data-Driven Ontology Validation
+Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology.
 Fig.
-3 shows that OCKGs comply with a star-based matching schema.
+3 shows that DOVA complies with a star-based matching schema.
 Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
-Figure 3: Ontology-compliant knowledge graphs (OCKGs).
+Figure 3: Data-driven ontology validation (DOVA).
+The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5].
 Building OCKGs is not easy.
-To determine which candidate ontology fits the KG(s) better, the KG needs to be modelled with each candidate.
-This process itself is essentially building an alignment of the candidate ontologies, informed by the KG under study.
+To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate.
+This process essentially involves aligning the candidate ontologies, informed by the KG under study.
 III An Ecosystem
@@
 4 illustrates an overview of an ecosystem for ontology interoperability.
-Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy.
-We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve interoperability.
+Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy.
+We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA) in different phases of the ontology engineering life cycle to achieve interoperability.
 Figure 4: An overview of the ecosystem.
@@
 5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
-While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown in the following.
-  1. (1)
+While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
+  1. (i)
     Observation Pattern (used in SOSA [7] and SSN [8]).
@@
 The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
-  2. (2)
+  2. (ii)
     Measurement Pattern (used in the early version of SAREF [9]).
@@
 The measurement is made by the sensor, which is a measure of the room related to the property temperature.
-  3. (3)
+  3. (iii)
     Event Pattern (used in DUL [10]). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
-  4. (4)
+  4. (iv)
     Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
+Figure 5: Modelling using different ODPs.
 On the other hand, differences in ontology characteristics, such as those listed below, need to be addressed in practice.
-  1. 1.
-    The definition of the concepts can be different.
-For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room.
+  1. (i)
+    The definitions of the concepts can differ.
+Temperature is defined as a property of the room in both the Observation Pattern and the Measurement Pattern, but in the Event Pattern it is treated as a quality measure of the room.
 Activity Pattern defines the temperature as a member of the room.
-  2. 2.
+  2. (ii)
     The levels of expressiveness are different.
-Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish whether it is a structural attachment (i.e.
-isPartOf) or only a spatial presence (i.e.
+Unlike Observation Pattern, Measurement Pattern, and Activity Pattern, which can only describe a sensor that is installed in the room, the Event Pattern can distinguish between a structural attachment (i.e.
+isPartOf) and a spatial presence (i.e.
 isLocatedIn).
-Figure 5: Modelling using different ODPs.
 IV-B Develop Phase
@@
 As illustrated in Fig.
-6, we can see that SOSA and SSN are very similar because they are built from the same ODPs.
+6, SOSA and SSN are very similar because they are built from the same ODPs.
 However, they are not exactly the same.
-The OM task is to find aligned concepts between two different ontologies.
+The OM task is to identify aligned concepts across two ontologies.
 SOSA and an updated SSN were co-developed, with a published alignment between them.
@@
 Figure 6: Aligning SOSA and SSN.
-  1. 1.
+  1. (i)
     Finding trivial alignments only considers the syntactic matching of the entity.
@@
 We also refer to these mappings as “seed mappings” because they serve as the basis for subsequent lexical and semantic matching.
-  2. 2.
+  2. (ii)
     Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity.
@@
 7 shows the OCKGs based on SOSA and SSN.
-We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts within the given KG.
-  1. 1.
+We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
+Figure 7: Build OCKGs using SOSA and SSN.
+  1. (i)
     The stimulus of the sensor “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
-  2. 2.
+  2. (ii)
     The actual feature of interest “Building 001”. The sensor observed the temperature of Room 001. It is not a feature of interest, but a sample of the temperature of Building 001.
-To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA.
-We also make several design choices between aligned mappings (retrieved from [13, 15, 16]).
-  1. 1.
+To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA, as well as make several design choices between aligned mappings (retrieved from [13, 15, 16]).
+  1. (i)
     We use ssn:Property to replace sosa:ObservableProperty because this class is more general and can be applied in a broader context for sensing, including actuation and sampling.
-  2. 2.
+  2. (ii)
     We use sosa:isHostedby to replace ssn:onPlatform because this property is more closely related to describing a sensor physically installed in the room.
-  3. 3.
+  3. (iii)
     We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest.
 Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
-Figure 7: Build OCKGs using SOSA and SSN.
 Fig.
 8 shows the updated ontology and the OCKG built on the updated ontology.
-Apparently, the updated ontology is more suitable for the provided KG instance.
+Apparently, the updated ontology is better suited to the provided KG instance.
 In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times.
@@
 Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks.
-As a complement to ODPs, OM&OV can align the concepts that are not captured by ODPs.
+As a complement to ODPs, OM&OV can align concepts that are not captured by ODPs.
 On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems.
 OM&OV is a critical middleware between ODPs and OCKGs.
-V Discussion & Conclusion
+V Contributions
+  1. (i)
+    ODPs: We systematically analyse four popular ontologies used in the smart building domain and observe that each building ontology has merits and demerits.
+We propose building ODPs to capture common and well-agreed concepts and relations in the building domain, including building spaces, equipment and systems, points and measurements [6].
+  2. (ii)
+    OM&OV: We develop a software system for OM and OV.
+Agent-OM [17] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner.
+Agent-OV [14] evolves from Agent-OM and explicitly targets OV tasks, including new task formulation and evaluation, with a cross-reference mechanism that compiles OM tasks.
+Agent-OM is used to create an LLM hallucination dataset [18] for the LLM leaderboard and LLM fine-tuning for OM.
+Agent-OM also participate in the Ontology Alignment Evaluation Initiative (OAEI) [19] 2025 campaign.
+  3. (iii)
+    DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance.
+Examples in the building domain showcase the efficiency of using OCKGs in ontology validation and decision-making [5].
+VI Discussion & Conclusion
 Ontology interoperability has been a long-standing issue studied in the Semantic Web community.
 Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings.
-Ontology-compliant knowledge graphs (OCKGs) aim to explore the interoperability between an ontology and its complementary KGs.
+Data-driven ontology validation aims to explore the interoperability between an ontology and its complementary KGs.
 While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability.
-We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and constructing OCKGs in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
+We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
 Supplemental Material Statement
@@
 The author thanks the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+VII Extended Literature Review
+Ontology interoperability is a fundamental but critical issue in sharing and reusing ontologies across different applications.
+While various techniques have been applied in this area, this survey provides a comprehensive review of three dominant state-of-the-art semantic web techniques for ontology interoperability: ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
+We systematically compare the pros and cons of each technique and identify opportunities to integrate them to build an ecosystem for better ontology interoperability.
+VII-A Motivation
+An ontology is defined as “a formal and explicit specification of a shared conceptualisation” [20].
+Since its inception in computer science, ontology has been applied in various applications to improve semantic interoperability for data integration and knowledge sharing.
+However, one fundamental but critical question is: Is the ontology itself interoperable?
+There is no doubt that using an ontology for knowledge graphs (KGs) can enhance KG interoperability through shared vocabularies and definitions, but the ontology itself is not interoperable.
+For example, Brick [21] and RECore [22] are two popular ontologies in the building domain, but their coverage is different and they do not share the same concepts.
+Brick offers more detailed concepts for building sensors and actuators, while RECore is better at presenting spatial information about the building.
+The two ontologies also use different terminologies to describe identical concepts.
+For example, the building’s physical location is referred to as ”Site” in Brick and ”Land” in RECore.
+A proposal [23] to harmonise these two ontologies began in 2022, but to date, many concepts have remained unaligned.
+Moreover, this collaboration not only aims to align two building ontologies but also includes many other emerging alignments, such as the Brick alignment to two industry standards, Industry Foundation Classes (IFC) [24] and ASHRAE Standard 223P [25].
+Ontology interoperability is an essential requirement for ontology engineering.
+It is not possible for two applications that use different ontologies to communicate without aligning their concepts.
+Several advanced topics, such as ontology modularisation [26] and ontology selection [27], are highly dependent on the level of ontology interoperability.
+When an ontology entity cannot find a replacement in another ontology, the extracted ontology module may treat it as missing, and this will lead to an entirely different ontology choice.
+In this paper, we provide a comprehensive review of state-of-the-art semantic web techniques for ontology interoperability.
+These include ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
+We review the current literature, identify knowledge gaps, and outline future directions.
+By leveraging their strengths while mitigating their weaknesses, we identify opportunities to build an ecosystem among these three techniques for better ontology interoperability.
+The rest of the paper is organised as follows.
+Section VII-B presents the preliminaries.
+Sections VII-CVII-D, and VII-E provide reviews of the current literature on ODPs, OM&OV, and DOVA, respectively.
+Section VII-F compares these three techniques and highlights their pros and cons in building an ecosystem.
+Section VII-G concludes the paper.
+VII-B Preliminaries
+VII-B1 Ontology
+The ontology is the cornerstone of knowledge graphs (KGs).
+KGs built from the ontology are much easier to interpret and understand.
+Fig.
+9 shows an example of using the ontology for KGs.
+The KG describes an event where Glen Powell and Sydney Sweeney travel to the city of Sydney to film a romantic comedy movie “Anyone But You” [1].
+The triples (Glen, travelsTo, Sydney) and (Sydney, travelsTo, Sydney) are snippets of the KG, while the triples (travelsTo, domain, Person) and (travelsTo, range, Place) are snippets of the ontology.
+The ontology snippets derived from the KG snippet (Glen, travelsTo, Sydney) can be used to explain the perplexing KG snippet (Sydney, travelsTo, Sydney).
+By using definitions provided by the ontology, it is easier to infer that the first subject entity “Sydney” refers to the person Sydney Sweeney, while the second object entity “Sydney” refers to the city of Sydney being visited.
+Figure 9: An example of using the ontology for explaining knowledge graphs.
+The ontology also provides extensive logical expressions for KGs.
+Fig.
+10 shows several examples of using logical expressions within the ontology.
+A person can travel to many places (Restriction 1), but can only live in one place (Restriction 2).
+A person can only travel to or live in a place (Restriction 3); the place can be a city, but it is not limited to the city alone (Restriction 4).
+These pieces of information cannot be represented in KGs, but can be modelled as machine-readable restrictions within the ontology.
+Figure 10: Several examples of using logical expressions within the ontology.
+Developing an ontology is not easy and requires extensive domain knowledge, but once it is developed, the ontology can be used across multiple KGs.
+Both the ontology and the KGs form a set of triples (subject, predicate, object).
+Still, ontology provides term definitions (e.g., class and property) as well as more advanced logical expressions (e.g.
+domain, range, and restriction).
+These terms can be modelled and formatted in the Web Ontology Language (OWL) [28] and other languages, such as Resource Description Framework Schema (RDFS) [29] and Simple Knowledge Organisation System (SKOS) [30].
+In general, the ontology serves as the conceptual model, whereas KGs store data instances.
+There is one exception: a data instance can be defined as an individual within the ontology, but its existence also means that the data instance (i.e.
+individual) is a well-agreed term that is not always changeable.
+For example, the term “Sydney” can be defined as an individual within the ontology.
+KGs do not always require an ontology.
+For some domains, ontology-free KGs are a more rational choice compared to ontology-based KGs.
+For example, building an ontology for social media is not recommended due to the rapid evolution of terms and the lack of universal agreement.
+VII-B2 Ontology Interoperability
+The ontology has been used for decades to improve the interoperability of KGs.
+However, a criticised view is that the ontology itself is not interoperable.
+An ontology can be built from scratch or reused from existing ontologies, but there is no guarantee that an ontology always contains the same concepts as other ontologies.
+Overlaps and missing concepts may exist, making it difficult to share and reuse an ontology in different applications.
+There are various types of interoperability issues in the ontology [31].
+  1. (i)
+    Formatting interoperability.
+The ontology can be formatted as OWL, RDF, or XML.
+There is a known issue with transferring ontologies to different formats due to differences in their levels of expressiveness.
+  2. (ii)
+    Terminological interoperability. The ontology can use different terminologies to describe the same concepts. For example, “AHU” and “Air Handling Unit”.
+  3. (iii)
+    Logical interoperability.
+The same concepts can have different meanings in different ontologies.
+For example, if the class “Meter” is the subclass of “Point”, it means a data point; if the class “Meter” is the subclass of “Device”, it means a physical device.
+  4. (iv)
+    Semantic interoperability.
+The same concepts can be modelled from different perspectives.
+For example, the class “HAVC Zone” can be logically considered as a functional unit or physically as a concrete spatial element.
+  5. (v)
+    Contextual interoperability.
+The same concepts can have different meanings in different domains.
+For example, the class “System” refers to a software application in the building domain, but means a set of organs in biology.
+VII-C Ontology Design Patterns
+VII-C1 Overview
+Ontology design patterns (ODPs) are built on the concept of modularisation and reusability.
+ODPs are the smallest and most reusable building blocks for constructing recurrent ontologies [32, 2, 33].
+ODPs are extracted or re-engineered from existing ontologies.
+For example, the Actuation-Actuator-Effect (AAE) ODP is extracted from SAN [34], while the Execution-Executor-Procedure (EEP) ODP is re-engineered from SEAS [35], SOSA [36], SSN [37], and DUL [38].
+Ontologies that follow the same ODP are interoperable.
+One can easily interpret identical classes, properties, and individuals based on the shared patterns.
+ODPs are not fully implemented as a real ontology and provide only high-level information on the design of the ontology.
+The fully implemented ontology is commonly referred to as the exemplary ontology, which is often confused with ODPs.
+Fig.
+11 shows an example of using the “part of” pattern in the ODP and the exemplary ontology for the building domain.
+The ODP contains only one “hasPart” and one “isPartOf”, but the exemplary ontology repeats these terms multiple times to model domain-specific part-whole relationships among buildings, floors, and rooms.
+Figure 11: Comparison of ontology design pattern and exemplary ontology.
+VII-C2 Ontology Design Patterns in the Wild
+ODPs have different types, including content ODPs, structural ODPs for logical and architectural issues, correspondence ODPs for reengineering and alignment, and other types such as reasoning, presentation, and lexico-syntactic ODPs.
+Content ODPs are the most widely used ODPs, defining patterns of classes, properties, and relations for domain ontologies [39].
+One of the most well-known ODP repositories is available at https://odpa.github.io/patterns-repository/ (formerly http://ontologydesignpatterns.org), with documentation on domain usage and competency questions.
+ODPs have been used in many domains, such as bioinformatics [40], geography [41], cultural heritage [42], material science [43], and energy efficiency in buildings [44].
+VII-C3 Knowledge Gaps
+  1. (i)
+    The content ODPs dominate the current ODPs. We do not see many other types of ODPs available, but they can be useful in some cases.
+  2. (ii)
+    Limited research is available on using machine learning (ML) or large language models (LLMs) to facilitate this process.
+The major problem is determining the level of abstraction for each ODP.
+Even with a number of samples provided, it seems impossible to apply a few-shot approach or train a model for this purpose.
+  3. (iii)
+    Current ODPs are handcrafted by humans and validated on a case-by-case basis.
+This makes it challenging to apply ODPs in complex and emerging domains.
+For example, the building domain encompasses a mixture of building spaces, systems, devices, sensors, and measurements [6].
+VII-D Ontology Matching and Versioning
+VII-D1 Overview
+Ontology matching (OM) aims to find similar concepts between different ontologies [3].
+OM is also known as ontology alignment (OA), but OA is more commonly used to describe the OM process.
+Ontology versioning (OV) aims to capture the differences between different versions of an ontology [4].
+In some cases, ontology evolution is a synonym of OV, but OV is preferred because the ontology is not evolved by nature but rather through human intervention.
+Ontologies are interoperable when their classes, properties, and individuals are explicitly aligned with OM&OV.
+One can easily transfer terms from one ontology to another using the provided reference alignment.
+Given a pair of ontologies, the OM&OV task can be formed by finding an alignment that contains a set of mappings [45].
+The most common OM&OV tasks target equivalence, as other relations (e.g.
+subsumption or disjointness) can be derived from equivalent mappings.
+Similarly, the OM&OV task can be more complex and extended to one-to-many and many-to-many mappings, but the most common OM&OV tasks target one-to-one mappings.
+VII-D2 Ontology Matching
+Fig.
+12 illustrates the roadmap of OM.
+Before ChatGPT (BC), a number of OM systems were developed using traditional logic-based expert systems and ML-based prediction systems.
+Following the introduction of ChatGPT in 2022 [46], there has been a brief period of using a purely LLM-based approach to perform OM tasks.
+Due to the current limitations of LLMs, the period after depression (AD) aims to use them with enhancements.
+Figure 12: Roadmap of ontology matching (OM).
+  1. (i)
+    Logic-based systems are highly dependent on domain expertise, and mapping is mainly based on the interpretation of logic expressions within the ontology.
+Typical logic-based OM systems include LogMap [47] [48], AgreementMakerLight (AML) [49] [50], and FCA-Map [51] [52].
+Such systems generally compile the results quickly with acceptable performance for simple OM tasks.
+However, these systems face a bottleneck that prevents them from performing complex OM tasks.
+It is almost impossible to capture a comprehensive list of logic expressions in various domains and applications.
+  2. (ii)
+    ML-based systems are derived from the popular concept of machine learning in the field of artificial intelligence.
+For example, LogMap-ML [53] and BERTMap [54].
+Such systems typically include a learning process that requires a certain amount of training data for the model to learn patterns and make predictions.
+As with ML applications in other domains, ML-based systems face the challenge of sampling high-quality training data from the full dataset.
+The explainability of model choice is a long-standing issue within the scope of Explainable Artificial Intelligence (XAI) [55].
+  3. (iii)
+    LLM-based systems refer to systems that use LLMs directly for OM.
+In the early stages of using LLMs for OM, a general approach is to feed ontologies into the LLM entirely with minimal instructions.
+Such examples can be found in [56] and [57].
+One fundamental issue is the scalability of this approach.
+For large-scale ontologies, it is not feasible to feed the entire ontology into LLMs, as these systems may exceed their input token limits.
+In addition, these systems often exhibit low precision and recall due to the inherent hallucinations associated with LLMs.
+  4. (iv)
+    LLM-enhanced systems hold the view that current LLMs are not perfect.
+It is required to re-engineer the LLM-based infrastructure for OM tasks.
+This field is a new research area with two main streams.
+One is using retrieval-augmented generation (RAG) [58] to inject external resources and furnish the knowledge base of LLMs.
+This often results in the introduction of novel LLM-based architectures for OM tasks, such as [17, 59, 60, 61, 62, 63].
+The other direction aims to feed external knowledge into LLMs by training or fine-tuning, as in [64, 18, 65, 66].
+Table I summarises the pros and cons of different types of OM systems.
+While LLM-enhanced architectures are the state-of-the-art solution for OM systems, they also suffer from high cost and non-determinism, which can be mitigated by integrating traditional system designs.
+TABLE I: Pros and cons of different types of ontology matching systems.
+System Type Pros Cons
+Logic-Based Acceptable performance for simple tasks Fast and requires low computational power Requires extensive domain expertise Meet the bottleneck for complex OM tasks
+ML-Based High accuracy in trained models Models can be generalised for different domains Lack of high-quality training data Lack of explainability
+LLM-Based Ease of use with intuitive prompt engineering Provides mapping explanations Not scalable for large-scale ontologies Low accuracy caused by LLM hallucinations
+LLM-Enhanced High accuracy in complex tasks Provides precise mapping explanations Extensive burden for system re-engineering Dynamic outputs due to the LLM non-determinism
+VII-D3 Ontology Versioning
+Fig.
+13 illustrates that there are two dominant approaches used in OV tasks.
+One approach is to capture versioning information within the ontology and expose it as additional triples.
+The other approach is to attach the versioning information outside the ontology, as a complement to the ontology.
+Figure 13: Roadmap of ontology versioning (OV).
+  1. (i)
+[…]
+  2. (ii)
+    There are many ways of attaching versioning information to the ontology.
+Versioning information can be captured as a log [69], a standalone KG [70], or a network of KGs [71].
+While this approach is more flexible in its support for formats and does not increase the size of the ontology, it requires additional control and stores versioning files.
+VII-D4 Knowledge Gaps
+  1. (i)
+    In the era of LLMs, LLMs offer extensive capabilities to perform OM tasks.
+It is questionable whether the logic and ML used in classical OM systems remain useful in modern LLM-based systems.
+A study in [72] shows that classical text preprocessing is not outdated and can serve as a supplement to modern LLM-based OM systems.
+  2. (ii)
+    LLMs are useful tools for OM, but OM systems are complex and lack centralised control, which cannot meet the requirements for large-scale automation.
+LLM agents could be an interesting area to explore [17].
+  3. (iii)
+    OV is similar to OM, but it is unclear whether the pipeline for OM tasks can be reused for OV tasks. The authors in [14] introduce a unified pipeline that combines OM and OV.
+TABLE II: Comparison of three state-of-the-art semantic web techniques for ontology interoperability: ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
+Technique Prerequisite Concept Coverage Alignment Approach Alignment Schema Medium Used
+ODPs Ontologies Only Commonalities Only Align Conceptual Model Hub-and-Spoke Common Patterns
+OM&OV Ontologies Only Commonalities + Differences Align Ontology Entities Pairwise N/A
+DOVA Ontologies + References Differences Only Align Data model Hub-and-Spoke Knowledge Graphs
+VII-E Data-Driven Ontology Validation
+VII-E1 Overview
+Ontology validation (OVA) aims to check the compliance of the ontology.
+OVA is sometimes confused with ontology verification (OVF), but the two have different goals.
+The former refers to “Do we build the ontology correctly?”, while the latter refers to “Do we build the correct ontology?”.
+OVF ensures that the ontology is intrinsically correct, while OVA ensures that the ontology is extrinsically meaningful.
+One common type of OVA is to apply ontologies to a specific task or application and validate them using data instances.
+This process is known as data-driven ontology validation (DOVA).
+Ontologies that conform to the same data are interoperable.
+One can easily determine mappings between ontologies if entities are assigned to the same instance.
+KGs are the data most commonly used for DOVA.
+Depending on different assumptions and baselines, DOVA can be categorised into two types: ontology-based KGs and ontology-aware KGs.
+Ontology-based KGs assume the ontology is complete, and KGs need to be modified if they do not conform to the ontology.
+On the other hand, ontology-aware KGs treat KGs as baselines, and the ontology needs to be modified if it does not conform to them.
+In the literature, ontology-aware KGs are also referred to as ontology enrichment [73] and ontology reshaping [74, 75, 76].
+VII-E2 Ontology-Based Knowledge Graphs
+Ontology-based KGs validate KGs based on the given ontology.
+For example, when the domain of ”Feature of Interest” is defined as observation and actuation in the ontology, linking this term to a sensor and actuator violates this rule.
+Shapes Constraint Language (SHACL) [77] is the W3C recommendation to validate KGs that comply with the ontology.
+SHACL provides a more comprehensive and extendable validation than the built-in logical expressions within the ontology.
+SHACL allows KGs to conform to a specific “shape” defined by the end user.
+This “shape” can be a simple triple, but it can also be very complex with a shape graph.
+VII-E3 Ontology-Aware Knowledge Graphs
+Ontology-aware KGs validate the ontology against the given KGs, identifying missing or unused concepts.
+Ontology enrichment is the process of adding missing concepts to the ontology.
+For example, if the KG contains ”Quality” under ”Feature of Interest”, then the ontology needs to add this new concept accordingly.
+Ontology reshaping is the process of removing unused concepts from the ontology.
+For example, if the KG is only related to sensor observation, then “Trigger” used for the actuator can be removed from the ontology.
+VII-E4 Knowledge Gaps
+  1. (i)
+    In reality, ontologies are inherently incomplete.
+Although ontology-based KGs have been applied to many domains, they are not always applicable, especially in emerging domains where concepts are evolving rapidly.
+  2. (ii)
+    While ontology enrichment specifies adding new concepts and ontology reshaping specifies deleting unused concepts, there is a need to support adding, deleting, and updating functions concurrently.
+Moreover, ontology-aware KGs assume that the selected ontology is the optimal choice for the given KGs.
+Limited attention has been paid to finding an optimal ontology from a number of ontology candidates or pattern-based ontology fragments [5].
+  3. (iii)
+    There are no standard evaluation metrics and datasets for DOVA.
+This limitation makes it difficult to choose an ontology.
+For example, the decision can vary when one concept is missing from one ontology and misinterpreted in another.
+VII-F Discussion
+Table II summarises the comparison of three state-of-the-art semantic web techniques for ontology interoperability.
+We can see that ODPs, OM&OV, and DOVA differ from several perspectives, including prerequisite, concept coverage, alignment approach and schema, and medium used.
+  1. (i)
+    ODPs provide high-level guidelines for building ontologies.
+ODPs achieve ontology interoperability by aligning the conceptual model.
+This approach often extracts common patterns across different ontologies in a hub and ignores their differences.
+ODPs can be reused across different ontologies.
+  2. (ii)
+    OM&OV aim to find the corresponding mappings between two ontologies.
+OM&OV achieve ontology interoperability by aligning ontology entities.
+While OM only finds aligned entities between ontologies (i.e.
+commonalities), OV also finds unaligned entities (i.e.
+differences).
+  3. (iii)
+    DOVA validates the ontologies from the data perspective.
+DOVA achieves ontology interoperability by aligning the data model.
+This approach typically applies a KG hub across multiple ontology spokes.
+The DOVA process uses reference alignments to identify commonalities between ontologies and thereby highlight their differences in modelling.
+These three techniques can work independently for ontology interoperability, but they also have some inter-dependencies.
+Therefore, there are opportunities to develop an ecosystem that leverages their strengths while mitigating their weaknesses.
+  1. (i)
+    ODPs provide only high-level guidelines for conceptual modelling and lack detailed mappings between classes, properties, and individuals.
+ODPs often need to cooperate with OM&OV and DOVA in downstream tasks.
+  2. (ii)
+    OM&OV provide detailed mappings for classes, properties, and individuals, but with limited support from the modelling perspective.
+ODPs can simplify the OM&OV process.
+Ontologies using the same ODP are much easier to align, as they form the same conceptual model.
+The output of OM&OV can serve as a reference for DOVA.
+  3. (iii)
+    DOVA is an evaluation process and does not provide any mappings by nature.
+DOVA reuses the commonalities generated by OM&OV and thereby identifies missing or overlapping modelling issues.
+The output of DOVA can be used to optimise existing ODPs in the next cycle.
+VII-G Conclusion
+While the induction of ontology aims to provide interoperability for KGs, with numerous ontologies developed for different domains and various purposes, ontology interoperability is becoming a bottleneck for sharing and reusing ontologies in real-world applications.
+In this paper, we review three dominant semantic techniques used for ontology interoperability.
+We observe that each technique targets a specific objective in ontology interoperability and differs in terms of strengths and weaknesses.
+The inter-dependencies between ODPs, OM&OV, and DOVA offer opportunities to develop an ecosystem for ontology interoperability.

### v6

@@
-An Ecosystem for Ontology Interoperability
+ISWC 2026 Companion Volume, October 25-29, 2026, Bari, Italy
 Abstract
@@
 A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
-I Introduction
-For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs.
+keywords:
+[…]
+1 Problem Statement
+For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability among KGs.
 While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs.
@@
 Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application.
-There is no guarantee that ontologies in the same domain agree on the conceptualisation.
-Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies.
-For example, it is common for ontologies to encompass a range of different concepts, such as describing a person using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee).
-Moreover, the terminology used in ontologies can be different.
-It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
-In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs).
-We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability.
-We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
-II Related Work
-II-A Ontology Design Patterns
-The goal of ontology design patterns (ODPs) is to identify common ontology fragments across different ontologies [2].
-Fig.
-1 illustrates that ODPs use a star-based matching schema.
-Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
-Figure 1: Ontology design patterns (ODPs).
-However, ODPs also have drawbacks.
+There is no guarantee that ontologies in the same domain share the same conceptualisation.
+Therefore, ontology interoperability has become a higher-level issue that affects the sharing and reuse of ontologies.
+Ontologies can encompass different entities to describe the same concepts.
+For example, describing a person can use their basic profiles (name, age, and gender), social relationships (parent-child and friendship), or professional roles (student and employee).
+The terminology used in the ontologies may also vary.
+One ontology can use “first name” and “last name” to describe a person’s full name, and another ontology can use “given name” and “family name” instead.
+To tackle this challenge, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA).
+We analyse their pros and cons and explore their roles in ontology engineering to improve interoperability.
+We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to demonstrate how these three techniques can foster strengths and circumvent weaknesses throughout the ontology engineering life cycle.
+2 Related Work
+2.1 Ontology Design Patterns (ODPs)
+Ontology design patterns (ODPs) identify common ontology fragments across different ontologies [2].
+ODPs use a star-based matching schema.
+Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies (see Figure 1).
+ODPs have some limitations.
 Firstly, ODPs are conceptual models and do not contain details of the actual class or property.
-This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices.
-Secondly, ODPs are commonly handcrafted on the basis of existing ontologies.
+This means that it is not possible to control terminology, and ontology creators remain free to make their own choices.
+Secondly, ODPs are commonly handcrafted based on existing ontologies.
 They are static and tend not to capture changes over time.
 Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
-II-B Ontology Matching and Versioning
+2.2 Ontology Matching and Versioning (OM&OV)
 Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4].
-Fig.
-2 shows that OM&OV conform to a pairwise matching schema.
-This means that each mapping is a pairwise process to find the correspondences between two ontologies.
-Figure 2: Ontology matching and versioning (OM&OV).
+OM&OV conform to a pairwise matching schema.
+This means that each mapping is a pairwise process to find the correspondences between two ontologies (see Figure 1).
 OM&OV are typically resource-intensive.
@@
 TBox-to-ABox matching).
-II-C Data-Driven Ontology Validation
+2.3 Data-Driven Ontology Validation (DOVA)
 Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology.
-Fig.
-3 shows that DOVA complies with a star-based matching schema.
-Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
-Figure 3: Data-driven ontology validation (DOVA).
+DOVA complies with a star-based matching schema.
+Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties that point to the same KG entities (see Figure 1).
 The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5].
 Building OCKGs is not easy.
-To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate.
+To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate ontology.
 This process essentially involves aligning the candidate ontologies, informed by the KG under study.
-III An Ecosystem
-Fig.
-4 illustrates an overview of an ecosystem for ontology interoperability.
+Figure 1: Illustration of different matching schema used in ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA).
+3 An Ecosystem for Ontology Interoperability
+Figure 2 illustrates an overview of the ecosystem for ontology interoperability.
 Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy.
-We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA) in different phases of the ontology engineering life cycle to achieve interoperability.
-Figure 4: An overview of the ecosystem.
-The design phase is responsible for understanding the requirements and creating a blueprint or prototype.
+We employ ODPs, OM&OV, and DOVA in different phases.
+(1) The design phase is responsible for understanding the requirements and creating a blueprint or prototype.
 In ontology engineering, these requirements are collected by competency questions.
 ODPs can serve as a skeleton to create conceptual models based on competency questions.
-The develop phase includes turning the prototype into a full product and applying a unit test.
+(2) The develop phase includes turning the prototype into a full product and applying a unit test.
 In ontology engineering, this phase is referred to as ontology construction and evaluation.
 OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version.
-The deploy phase usually refers to implementing the product in the real world.
+(3) The deploy phase usually refers to implementing the product in the real world.
 In ontology engineering, we refer to this as applying the ontology to actual KGs.
-We can utilise OCKGs to comprehend the interoperability of ontologies within each KG or a group of KGs, thereby enhancing the design of the ontology in the next cycle.
-IV A Case Study in the Building Domain
+We can utilise OCKGs to assess the interoperability of ontologies within a single KG or across a set of KGs, thereby enhancing the design of the ontology in the next cycle.
+Figure 2: An overview of the ecosystem. We propose to use ODPs in the design phase, OM&OV in the develop phase, and DOVA in the deploy phase to achieve better ontology interoperability.
+4 A Case Study in the Building Domain
 Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades.
@@
 The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
-IV-A Design Phase
+4.1 Ontology Design Patterns (ODPs) in the Design Phase
 Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability.
 It is commonly recommended to reuse pre-existing ODPs.
-Fig.
-5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
-While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
+Figure 3 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
   1. (i)
@@
     Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
-Figure 5: Modelling using different ODPs.
-On the other hand, differences in ontology characteristics, such as those listed below, need to be addressed in practice.
+While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
   1. (i)
@@
 isLocatedIn).
-IV-B Develop Phase
+Figure 3: Modelling sensor observation using different ODPs.
+4.2 Ontology Matching and Versioning (OM&OV) in the Develop Phase
 We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]).
-As illustrated in Fig.
-6, SOSA and SSN are very similar because they are built from the same ODPs.
+As illustrated in Figure 4, SOSA and SSN are very similar because they are built from the same ODPs.
 However, they are not exactly the same.
@@
 For example, ssn:hasProperty should not have a direct equivalent mapping in SOSA.
-IV-C Deploy Phase
+4.3 Data-Driven Ontology Validation (DOVA) in the Deploy Phase
 Now we have developed SOSA and SSN, along with their reference alignment.
@@
 Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation.
-Fig.
-7 shows the OCKGs based on SOSA and SSN.
+Figure 5 shows the OCKGs based on SOSA and SSN.
 We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
@@
     We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest.
-Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
-Fig.
-8 shows the updated ontology and the OCKG built on the updated ontology.
+Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we add the verb “has” to distinguish the property from the class.
+Figure 6 shows the updated ontology and the OCKG built on the updated ontology.
 Apparently, the updated ontology is better suited to the provided KG instance.
@@
 Figure 6: Updated ontology and new OCKG.
-IV-D Implications
+4.4 Reflection and Implications
 Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks.
@@
 OM&OV is a critical middleware between ODPs and OCKGs.
-V Contributions
+5 Contributions Made So Far
   1. (i)
@@
   2. (ii)
-    OM&OV: We develop a software system for OM and OV.
+    OM&OV: We study the interlink between modern LLMs and traditional OM systems [17].
+We develop a novel software system for OM and OV.
 Agent-OM [18] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner.
@@
 Agent-OM is used to create an LLM hallucination dataset [19] for the LLM leaderboard and LLM fine-tuning for OM.
-Agent-OM also participate in the Ontology Alignment Evaluation Initiative (OAEI) [19] 2025 campaign.
+Agent-OM also participates in the Ontology Alignment Evaluation Initiative (OAEI) [20] 2025 campaign and showcases its generative capability in handling OM tasks at different levels of granularities [21].
   3. (iii)
     DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance.
-Examples in the building domain showcase the efficiency of using OCKGs in ontology validation and decision-making [5].
-VI Discussion & Conclusion
+Examples in the building domain demonstrate the process of constraining OCKGs and showcase the efficiency of OCKGs in facilitating ontology validation and decision-making [5].
+6 Conclusion and Future Work
 Ontology interoperability has been a long-standing issue studied in the Semantic Web community.
 Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings.
-Data-driven ontology validation aims to explore the interoperability between an ontology and its complementary KGs.
+Data-driven ontology validation, which focuses on exploring the interoperability between an ontology and its complementary KGs, has received increasing attention in recent years.
 While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability.
-We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
-Supplemental Material Statement
-The source code, data, and/or other artifacts used in this study are available at https://github.com/qzc438-research/ontology-interoperability.
-Acknowledgment
-The author is supervised by Kerry Taylor (Australian National University), Weiqing Wang (Monash University), Jing Jiang (Australian National University), Subbu Sethuvenkatraman (Commonwealth Scientific and Industrial Research Organisation), Qing Wang (Australian National University), and Alex Potanin (Australian National University).
-The author thanks the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
-VII Extended Literature Review
-Ontology interoperability is a fundamental but critical issue in sharing and reusing ontologies across different applications.
-While various techniques have been applied in this area, this survey provides a comprehensive review of three dominant state-of-the-art semantic web techniques for ontology interoperability: ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
-We systematically compare the pros and cons of each technique and identify opportunities to integrate them to build an ecosystem for better ontology interoperability.
-VII-A Motivation
-An ontology is defined as “a formal and explicit specification of a shared conceptualisation” [20].
-Since its inception in computer science, ontology has been applied in various applications to improve semantic interoperability for data integration and knowledge sharing.
-However, one fundamental but critical question is: Is the ontology itself interoperable?
-There is no doubt that using an ontology for knowledge graphs (KGs) can enhance KG interoperability through shared vocabularies and definitions, but the ontology itself is not interoperable.
-For example, Brick [21] and RECore [22] are two popular ontologies in the building domain, but their coverage is different and they do not share the same concepts.
-Brick offers more detailed concepts for building sensors and actuators, while RECore is better at presenting spatial information about the building.
-The two ontologies also use different terminologies to describe identical concepts.
-For example, the building’s physical location is referred to as ”Site” in Brick and ”Land” in RECore.
-A proposal [23] to harmonise these two ontologies began in 2022, but to date, many concepts have remained unaligned.
-Moreover, this collaboration not only aims to align two building ontologies but also includes many other emerging alignments, such as the Brick alignment to two industry standards, Industry Foundation Classes (IFC) [24] and ASHRAE Standard 223P [25].
-Ontology interoperability is an essential requirement for ontology engineering.
-It is not possible for two applications that use different ontologies to communicate without aligning their concepts.
-Several advanced topics, such as ontology modularisation [26] and ontology selection [27], are highly dependent on the level of ontology interoperability.
-When an ontology entity cannot find a replacement in another ontology, the extracted ontology module may treat it as missing, and this will lead to an entirely different ontology choice.
-In this paper, we provide a comprehensive review of state-of-the-art semantic web techniques for ontology interoperability.
-These include ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
-We review the current literature, identify knowledge gaps, and outline future directions.
-By leveraging their strengths while mitigating their weaknesses, we identify opportunities to build an ecosystem among these three techniques for better ontology interoperability.
-The rest of the paper is organised as follows.
-Section VII-B presents the preliminaries.
-Sections VII-CVII-D, and VII-E provide reviews of the current literature on ODPs, OM&OV, and DOVA, respectively.
-Section VII-F compares these three techniques and highlights their pros and cons in building an ecosystem.
-Section VII-G concludes the paper.
-VII-B Preliminaries
-VII-B1 Ontology
-The ontology is the cornerstone of knowledge graphs (KGs).
-KGs built from the ontology are much easier to interpret and understand.
-Fig.
-9 shows an example of using the ontology for KGs.
-The KG describes an event where Glen Powell and Sydney Sweeney travel to the city of Sydney to film a romantic comedy movie “Anyone But You” [1].
-The triples (Glen, travelsTo, Sydney) and (Sydney, travelsTo, Sydney) are snippets of the KG, while the triples (travelsTo, domain, Person) and (travelsTo, range, Place) are snippets of the ontology.
-The ontology snippets derived from the KG snippet (Glen, travelsTo, Sydney) can be used to explain the perplexing KG snippet (Sydney, travelsTo, Sydney).
-By using definitions provided by the ontology, it is easier to infer that the first subject entity “Sydney” refers to the person Sydney Sweeney, while the second object entity “Sydney” refers to the city of Sydney being visited.
-Figure 9: An example of using the ontology for explaining knowledge graphs.
-The ontology also provides extensive logical expressions for KGs.
-Fig.
-10 shows several examples of using logical expressions within the ontology.
-A person can travel to many places (Restriction 1), but can only live in one place (Restriction 2).
-A person can only travel to or live in a place (Restriction 3); the place can be a city, but it is not limited to the city alone (Restriction 4).
-These pieces of information cannot be represented in KGs, but can be modelled as machine-readable restrictions within the ontology.
-Figure 10: Several examples of using logical expressions within the ontology.
-Developing an ontology is not easy and requires extensive domain knowledge, but once it is developed, the ontology can be used across multiple KGs.
-Both the ontology and the KGs form a set of triples (subject, predicate, object).
-Still, ontology provides term definitions (e.g., class and property) as well as more advanced logical expressions (e.g.
-domain, range, and restriction).
-These terms can be modelled and formatted in the Web Ontology Language (OWL) [28] and other languages, such as Resource Description Framework Schema (RDFS) [29] and Simple Knowledge Organisation System (SKOS) [30].
-In general, the ontology serves as the conceptual model, whereas KGs store data instances.
-There is one exception: a data instance can be defined as an individual within the ontology, but its existence also means that the data instance (i.e.
-individual) is a well-agreed term that is not always changeable.
-For example, the term “Sydney” can be defined as an individual within the ontology.
-KGs do not always require an ontology.
-For some domains, ontology-free KGs are a more rational choice compared to ontology-based KGs.
-For example, building an ontology for social media is not recommended due to the rapid evolution of terms and the lack of universal agreement.
-VII-B2 Ontology Interoperability
-The ontology has been used for decades to improve the interoperability of KGs.
-However, a criticised view is that the ontology itself is not interoperable.
-An ontology can be built from scratch or reused from existing ontologies, but there is no guarantee that an ontology always contains the same concepts as other ontologies.
-Overlaps and missing concepts may exist, making it difficult to share and reuse an ontology in different applications.
-There are various types of interoperability issues in the ontology [31].
-  1. (i)
-    Formatting interoperability.
-The ontology can be formatted as OWL, RDF, or XML.
-There is a known issue with transferring ontologies to different formats due to differences in their levels of expressiveness.
-  2. (ii)
-    Terminological interoperability. The ontology can use different terminologies to describe the same concepts. For example, “AHU” and “Air Handling Unit”.
-  3. (iii)
-    Logical interoperability.
-The same concepts can have different meanings in different ontologies.
-For example, if the class “Meter” is the subclass of “Point”, it means a data point; if the class “Meter” is the subclass of “Device”, it means a physical device.
-  4. (iv)
-    Semantic interoperability.
-The same concepts can be modelled from different perspectives.
-For example, the class “HAVC Zone” can be logically considered as a functional unit or physically as a concrete spatial element.
-  5. (v)
-    Contextual interoperability.
-The same concepts can have different meanings in different domains.
-For example, the class “System” refers to a software application in the building domain, but means a set of organs in biology.
-VII-C Ontology Design Patterns
-VII-C1 Overview
-Ontology design patterns (ODPs) are built on the concept of modularisation and reusability.
-ODPs are the smallest and most reusable building blocks for constructing recurrent ontologies [32, 2, 33].
-ODPs are extracted or re-engineered from existing ontologies.
-For example, the Actuation-Actuator-Effect (AAE) ODP is extracted from SAN [34], while the Execution-Executor-Procedure (EEP) ODP is re-engineered from SEAS [35], SOSA [36], SSN [37], and DUL [38].
-Ontologies that follow the same ODP are interoperable.
-One can easily interpret identical classes, properties, and individuals based on the shared patterns.
-ODPs are not fully implemented as a real ontology and provide only high-level information on the design of the ontology.
-The fully implemented ontology is commonly referred to as the exemplary ontology, which is often confused with ODPs.
-Fig.
-11 shows an example of using the “part of” pattern in the ODP and the exemplary ontology for the building domain.
-The ODP contains only one “hasPart” and one “isPartOf”, but the exemplary ontology repeats these terms multiple times to model domain-specific part-whole relationships among buildings, floors, and rooms.
-Figure 11: Comparison of ontology design pattern and exemplary ontology.
-VII-C2 Ontology Design Patterns in the Wild
-ODPs have different types, including content ODPs, structural ODPs for logical and architectural issues, correspondence ODPs for reengineering and alignment, and other types such as reasoning, presentation, and lexico-syntactic ODPs.
-Content ODPs are the most widely used ODPs, defining patterns of classes, properties, and relations for domain ontologies [39].
-One of the most well-known ODP repositories is available at https://odpa.github.io/patterns-repository/ (formerly http://ontologydesignpatterns.org), with documentation on domain usage and competency questions.
-ODPs have been used in many domains, such as bioinformatics [40], geography [41], cultural heritage [42], material science [43], and energy efficiency in buildings [44].
-VII-C3 Knowledge Gaps
-  1. (i)
-    The content ODPs dominate the current ODPs. We do not see many other types of ODPs available, but they can be useful in some cases.
-  2. (ii)
-    Limited research is available on using machine learning (ML) or large language models (LLMs) to facilitate this process.
-The major problem is determining the level of abstraction for each ODP.
-Even with a number of samples provided, it seems impossible to apply a few-shot approach or train a model for this purpose.
-  3. (iii)
-    Current ODPs are handcrafted by humans and validated on a case-by-case basis.
-This makes it challenging to apply ODPs in complex and emerging domains.
-For example, the building domain encompasses a mixture of building spaces, systems, devices, sensors, and measurements [6].
-VII-D Ontology Matching and Versioning
-VII-D1 Overview
-Ontology matching (OM) aims to find similar concepts between different ontologies [3].
-OM is also known as ontology alignment (OA), but OA is more commonly used to describe the OM process.
-Ontology versioning (OV) aims to capture the differences between different versions of an ontology [4].
-In some cases, ontology evolution is a synonym of OV, but OV is preferred because the ontology is not evolved by nature but rather through human intervention.
-Ontologies are interoperable when their classes, properties, and individuals are explicitly aligned with OM&OV.
-One can easily transfer terms from one ontology to another using the provided reference alignment.
-Given a pair of ontologies, the OM&OV task can be formed by finding an alignment that contains a set of mappings [45].
-The most common OM&OV tasks target equivalence, as other relations (e.g.
-subsumption or disjointness) can be derived from equivalent mappings.
-Similarly, the OM&OV task can be more complex and extended to one-to-many and many-to-many mappings, but the most common OM&OV tasks target one-to-one mappings.
-VII-D2 Ontology Matching
-Fig.
-12 illustrates the roadmap of OM.
-Before ChatGPT (BC), a number of OM systems were developed using traditional logic-based expert systems and ML-based prediction systems.
-Following the introduction of ChatGPT in 2022 [46], there has been a brief period of using a purely LLM-based approach to perform OM tasks.
-Due to the current limitations of LLMs, the period after depression (AD) aims to use them with enhancements.
-Figure 12: Roadmap of ontology matching (OM).
-  1. (i)
-    Logic-based systems are highly dependent on domain expertise, and mapping is mainly based on the interpretation of logic expressions within the ontology.
-Typical logic-based OM systems include LogMap [47] [48], AgreementMakerLight (AML) [49] [50], and FCA-Map [51] [52].
-Such systems generally compile the results quickly with acceptable performance for simple OM tasks.
-However, these systems face a bottleneck that prevents them from performing complex OM tasks.
-It is almost impossible to capture a comprehensive list of logic expressions in various domains and applications.
-  2. (ii)
-    ML-based systems are derived from the popular concept of machine learning in the field of artificial intelligence.
-For example, LogMap-ML [53] and BERTMap [54].
-Such systems typically include a learning process that requires a certain amount of training data for the model to learn patterns and make predictions.
-As with ML applications in other domains, ML-based systems face the challenge of sampling high-quality training data from the full dataset.
-The explainability of model choice is a long-standing issue within the scope of Explainable Artificial Intelligence (XAI) [55].
-  3. (iii)
-    LLM-based systems refer to systems that use LLMs directly for OM.
-In the early stages of using LLMs for OM, a general approach is to feed ontologies into the LLM entirely with minimal instructions.
-Such examples can be found in [56] and [57].
-One fundamental issue is the scalability of this approach.
-For large-scale ontologies, it is not feasible to feed the entire ontology into LLMs, as these systems may exceed their input token limits.
-In addition, these systems often exhibit low precision and recall due to the inherent hallucinations associated with LLMs.
-  4. (iv)
-    LLM-enhanced systems hold the view that current LLMs are not perfect.
-It is required to re-engineer the LLM-based infrastructure for OM tasks.
-This field is a new research area with two main streams.
-One is using retrieval-augmented generation (RAG) [58] to inject external resources and furnish the knowledge base of LLMs.
-This often results in the introduction of novel LLM-based architectures for OM tasks, such as [17, 59, 60, 61, 62, 63].
-The other direction aims to feed external knowledge into LLMs by training or fine-tuning, as in [64, 18, 65, 66].
-Table I summarises the pros and cons of different types of OM systems.
-While LLM-enhanced architectures are the state-of-the-art solution for OM systems, they also suffer from high cost and non-determinism, which can be mitigated by integrating traditional system designs.
-TABLE I: Pros and cons of different types of ontology matching systems.
-System Type Pros Cons
-Logic-Based Acceptable performance for simple tasks Fast and requires low computational power Requires extensive domain expertise Meet the bottleneck for complex OM tasks
-ML-Based High accuracy in trained models Models can be generalised for different domains Lack of high-quality training data Lack of explainability
-LLM-Based Ease of use with intuitive prompt engineering Provides mapping explanations Not scalable for large-scale ontologies Low accuracy caused by LLM hallucinations
-LLM-Enhanced High accuracy in complex tasks Provides precise mapping explanations Extensive burden for system re-engineering Dynamic outputs due to the LLM non-determinism
-VII-D3 Ontology Versioning
-Fig.
-13 illustrates that there are two dominant approaches used in OV tasks.
-One approach is to capture versioning information within the ontology and expose it as additional triples.
-The other approach is to attach the versioning information outside the ontology, as a complement to the ontology.
-Figure 13: Roadmap of ontology versioning (OV).
-  1. (i)
-[…]
-  2. (ii)
-    There are many ways of attaching versioning information to the ontology.
-Versioning information can be captured as a log [69], a standalone KG [70], or a network of KGs [71].
-While this approach is more flexible in its support for formats and does not increase the size of the ontology, it requires additional control and stores versioning files.
-VII-D4 Knowledge Gaps
-  1. (i)
-    In the era of LLMs, LLMs offer extensive capabilities to perform OM tasks.
-It is questionable whether the logic and ML used in classical OM systems remain useful in modern LLM-based systems.
-A study in [72] shows that classical text preprocessing is not outdated and can serve as a supplement to modern LLM-based OM systems.
-  2. (ii)
-    LLMs are useful tools for OM, but OM systems are complex and lack centralised control, which cannot meet the requirements for large-scale automation.
-LLM agents could be an interesting area to explore [17].
-  3. (iii)
-    OV is similar to OM, but it is unclear whether the pipeline for OM tasks can be reused for OV tasks. The authors in [14] introduce a unified pipeline that combines OM and OV.
-TABLE II: Comparison of three state-of-the-art semantic web techniques for ontology interoperability: ontology design patterns (ODPs), ontology matching and ontology versioning (OM&OV), and data-driven ontology validation (DOVA).
-Technique Prerequisite Concept Coverage Alignment Approach Alignment Schema Medium Used
-ODPs Ontologies Only Commonalities Only Align Conceptual Model Hub-and-Spoke Common Patterns
-OM&OV Ontologies Only Commonalities + Differences Align Ontology Entities Pairwise N/A
-DOVA Ontologies + References Differences Only Align Data model Hub-and-Spoke Knowledge Graphs
-VII-E Data-Driven Ontology Validation
-VII-E1 Overview
-Ontology validation (OVA) aims to check the compliance of the ontology.
-OVA is sometimes confused with ontology verification (OVF), but the two have different goals.
-The former refers to “Do we build the ontology correctly?”, while the latter refers to “Do we build the correct ontology?”.
-OVF ensures that the ontology is intrinsically correct, while OVA ensures that the ontology is extrinsically meaningful.
-One common type of OVA is to apply ontologies to a specific task or application and validate them using data instances.
-This process is known as data-driven ontology validation (DOVA).
-Ontologies that conform to the same data are interoperable.
-One can easily determine mappings between ontologies if entities are assigned to the same instance.
-KGs are the data most commonly used for DOVA.
-Depending on different assumptions and baselines, DOVA can be categorised into two types: ontology-based KGs and ontology-aware KGs.
-Ontology-based KGs assume the ontology is complete, and KGs need to be modified if they do not conform to the ontology.
-On the other hand, ontology-aware KGs treat KGs as baselines, and the ontology needs to be modified if it does not conform to them.
-In the literature, ontology-aware KGs are also referred to as ontology enrichment [73] and ontology reshaping [74, 75, 76].
-VII-E2 Ontology-Based Knowledge Graphs
-Ontology-based KGs validate KGs based on the given ontology.
-For example, when the domain of ”Feature of Interest” is defined as observation and actuation in the ontology, linking this term to a sensor and actuator violates this rule.
-Shapes Constraint Language (SHACL) [77] is the W3C recommendation to validate KGs that comply with the ontology.
-SHACL provides a more comprehensive and extendable validation than the built-in logical expressions within the ontology.
-SHACL allows KGs to conform to a specific “shape” defined by the end user.
-This “shape” can be a simple triple, but it can also be very complex with a shape graph.
-VII-E3 Ontology-Aware Knowledge Graphs
-Ontology-aware KGs validate the ontology against the given KGs, identifying missing or unused concepts.
-Ontology enrichment is the process of adding missing concepts to the ontology.
-For example, if the KG contains ”Quality” under ”Feature of Interest”, then the ontology needs to add this new concept accordingly.
-Ontology reshaping is the process of removing unused concepts from the ontology.
-For example, if the KG is only related to sensor observation, then “Trigger” used for the actuator can be removed from the ontology.
-VII-E4 Knowledge Gaps
-  1. (i)
-    In reality, ontologies are inherently incomplete.
-Although ontology-based KGs have been applied to many domains, they are not always applicable, especially in emerging domains where concepts are evolving rapidly.
-  2. (ii)
-    While ontology enrichment specifies adding new concepts and ontology reshaping specifies deleting unused concepts, there is a need to support adding, deleting, and updating functions concurrently.
-Moreover, ontology-aware KGs assume that the selected ontology is the optimal choice for the given KGs.
-Limited attention has been paid to finding an optimal ontology from a number of ontology candidates or pattern-based ontology fragments [5].
-  3. (iii)
-    There are no standard evaluation metrics and datasets for DOVA.
-This limitation makes it difficult to choose an ontology.
-For example, the decision can vary when one concept is missing from one ontology and misinterpreted in another.
-VII-F Discussion
-Table II summarises the comparison of three state-of-the-art semantic web techniques for ontology interoperability.
-We can see that ODPs, OM&OV, and DOVA differ from several perspectives, including prerequisite, concept coverage, alignment approach and schema, and medium used.
-  1. (i)
-    ODPs provide high-level guidelines for building ontologies.
-ODPs achieve ontology interoperability by aligning the conceptual model.
-This approach often extracts common patterns across different ontologies in a hub and ignores their differences.
-ODPs can be reused across different ontologies.
-  2. (ii)
-    OM&OV aim to find the corresponding mappings between two ontologies.
-OM&OV achieve ontology interoperability by aligning ontology entities.
-While OM only finds aligned entities between ontologies (i.e.
-commonalities), OV also finds unaligned entities (i.e.
-differences).
-  3. (iii)
-    DOVA validates the ontologies from the data perspective.
-DOVA achieves ontology interoperability by aligning the data model.
-This approach typically applies a KG hub across multiple ontology spokes.
-The DOVA process uses reference alignments to identify commonalities between ontologies and thereby highlight their differences in modelling.
-These three techniques can work independently for ontology interoperability, but they also have some inter-dependencies.
-Therefore, there are opportunities to develop an ecosystem that leverages their strengths while mitigating their weaknesses.
-  1. (i)
-    ODPs provide only high-level guidelines for conceptual modelling and lack detailed mappings between classes, properties, and individuals.
-ODPs often need to cooperate with OM&OV and DOVA in downstream tasks.
-  2. (ii)
-    OM&OV provide detailed mappings for classes, properties, and individuals, but with limited support from the modelling perspective.
-ODPs can simplify the OM&OV process.
-Ontologies using the same ODP are much easier to align, as they form the same conceptual model.
-The output of OM&OV can serve as a reference for DOVA.
-  3. (iii)
-    DOVA is an evaluation process and does not provide any mappings by nature.
-DOVA reuses the commonalities generated by OM&OV and thereby identifies missing or overlapping modelling issues.
-The output of DOVA can be used to optimise existing ODPs in the next cycle.
-VII-G Conclusion
-While the induction of ontology aims to provide interoperability for KGs, with numerous ontologies developed for different domains and various purposes, ontology interoperability is becoming a bottleneck for sharing and reusing ontologies in real-world applications.
-In this paper, we review three dominant semantic techniques used for ontology interoperability.
-We observe that each technique targets a specific objective in ontology interoperability and differs in terms of strengths and weaknesses.
-The inter-dependencies between ODPs, OM&OV, and DOVA offer opportunities to develop an ecosystem for ontology interoperability.
+We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve ontology interoperability throughout the ontology engineering life cycle.
+Acknowledgements.
+The author is supervised by Kerry Taylor (Australian National University, ANU), Weiqing Wang (Monash University), Jing Jiang (ANU), Subbu Sethuvenkatraman (Commonwealth Scientific and Industrial Research Organisation, CSIRO), Qing Wang (ANU), and Alex Potanin (ANU).
+The author thanks CSIRO for supporting this project.
+Declaration on Generative AI
+During the preparation of this work, the authors used Grammarly in order to grammar and spell check, and to improve the text readability.
+After using the tool, the authors reviewed and edited the content and take full responsibility for the publication’s content.